4√2048*3√32=
2√320+4√80=
19√75-3√48=
21√27-3√75=
4√3072+4√75=
Can you also show the work?
1
Expert's answer
2012-03-20T09:52:48-0400
4√2048*3√32= 3*4*sqrt(2048*32)=12sqrt(65536)=12*256=3072

2√320+4√80==2*sqrt(4*80)+4sqty(80)=8sqty(80)

19√75-3√48=19sqrt(25*3)-3sqrt(16*3)=5*19sqrt(3)-3*4sqrt(3)=(95-12)sqrt(3)=83sqrt(3)
and so on